### v3

# The Quasar ULAS J1120+0641 May Never Reach a āNormalā Black Hole to Stellar Mass Ratio

Abstract
JWST observations of quasars in the Epoch of Reionization have revealed that many lie in host galaxies that are severely undermassive relative to the supermassive black holes. It is unclear how these systems will evolve to the tight local relation between stellar mass and black hole mass. We search for companions around the quasar ULAS J1120+0641 using JWST/NIRCam narrow, medium, and wide-band photometry to identify [OĀ III] emitters at the quasar redshift, and explore the potential for growth of the host galaxy through future mergers. We find 22 sources near the quasarās redshift across our two 4.4 arcmin2 fields, indicating that environment of ULAS J1120+0641 is strongly overdense in galaxies relative to the field. We estimate the potential future mass budget of the quasar host galaxy by summing the current stellar and gas masses of the quasar host and surrounding galaxies, correcting for incompleteness and selection effects. With no further black hole growth, ULAS J1120+0641 is unlikely to reach a ratio less than at , still much higher than typical for local galaxies. However, such systemsāa quiescent black hole in a low-luminosity galaxyāmay have escaped detection locally if they are sufficiently distant.
1 Introduction
Bright quasarsāthe sites of massive, fast-growing supermassive black holes (SMBHs)āprovide a fascinating laboratory to study the coevolution of SMBHs and their host galaxies, the physics of gas transport from the largest to smallest scales, and the effects of SMBH feedback on galaxies. The extreme luminosities of quasars have allowed their redshift frontier to expand swiftly from the local Universe (Schmidt, 1963) back to the Epoch of Reionization at (BogdĆ”n etĀ al., 2024; Fan etĀ al., 2001; Mortlock etĀ al., 2011; Fan etĀ al., 2023). However, the study of quasar hosts (outshined by the bright quasar point source) and the much fainter galaxies in their environments is significantly more difficult, especially at high redshift. Both areas of research have expanded significantly in recent years with JWST.
Quasars must accrete to produce their luminosity, indicating they are likely to live in gas-rich environments (at least, in the black holeās immediate neighborhood). On larger scales, though, recent JWST results have indicated that some luminous quasars at live in undermassive host galaxies compared to the local Universe (Onoue etĀ al., 2024; Stone etĀ al., 2024; Yue etĀ al., 2024), though with some exceptions (Ding etĀ al., 2023). Unfortunately, the difficulty of measuring the stellar mass of a bright quasarās host galaxy, especially at high redshift, means only a small number of high-redshift quasars currently have robust stellar masses. With this small sample, we cannot yet determine whether the behavior of these quasars reflects a genuine change in the relationship between black hole mass and stellar mass () at early times, or is simply a consequence of selection biases (Sun etĀ al., 2025b).
However, the fact remains that such extreme galaxies, with , are not observed in the local Universe aside from a few niche cases where the host galaxy has lost significant mass to tidal stripping. If these quasar host galaxies are to grow enough to land on the observed local relationship between SMBH mass and galaxy mass by , they must be able to form stars and merge with other galaxies in their environment. This fact, combined with predictions of the CDM paradigm and results of cosmological simulations, means when we observe quasars at the end of the Epoch of Reionization, we expect to see them living in dense environments, surrounded by other galaxies (Overzier etĀ al., 2009a; Inayoshi etĀ al., 2020; Volonteri etĀ al., 2021).
For decades, observations have targeted high-redshift quasars in hopes of detecting galaxy overdensities surrounding them. At high redshift (), HST and ground-based optical/NIR telescopes can only access the rest-frame ultraviolet, and must rely on the Lyman break technique (Adams etĀ al., 2015; Husband etĀ al., 2013; Morselli etĀ al., 2014; Ota etĀ al., 2018; Pudoka etĀ al., 2024; Stiavelli etĀ al., 2005; Zheng etĀ al., 2006) or selection of Lyman-alpha emitters (Chanchaiworawit etĀ al., 2019; Harikane etĀ al., 2019; Hennawi etĀ al., 2015; Kashikawa etĀ al., 2007; Overzier etĀ al., 2009b), with relatively few bands longward of the Lyman break, to identify candidate galaxies. Although galaxies were detected around many quasars, it remained difficult to characterize the environments of high-redshift quasars without access to the rest-frame optical at high sensitivity and resolution.
The study of high-redshift quasar environments was advanced significantly by the launch of JWST, which provides photometric bands extending much further longward of the Lyman break, and spectroscopy able to probe rest-optical emission lines, rather than Lyman- alone, to spectroscopically confirm overdensities surrounding high-redshift quasars (Champagne et al., 2025a; Kashino et al., 2023; Wang et al., 2023) and Little Red Dots (Schindler et al., 2024). While many studies with HST, ground-based telescopes, and now JWST have detected significant overdensities around targeted quasars, which may evolve into the massive galaxy clusters observed in the low-redshift Universe, some quasars do not have significant surrounding overdensities (see e.g. Bañados et al., 2013; Eilers et al., 2024; Rojas-Ruiz et al., 2024; Pudoka et al., 2025), and instead live in underdense environments, or environments consistent with the field.
A particularly interesting example of a quasar apparently without a surrounding overdensity is the quasar ULAS J112001.48+064124.3 (hereafter J1120+0641). Even among the (already extreme) population of bright high-redshift quasars, J1120+0641 is an outlier. At , it was at the time of its discovery the highest-redshift quasar known (Mortlock etĀ al., 2011), and measurements from single-epoch spectroscopy of various UV, optical, and near-IR lines place its black hole mass somewhere around Mā, only about 750 Myr after the Big Bang. Its host galaxy is extremely bright in the far-infrared (Venemans etĀ al., 2020), indicating significant dust-obscured star formation; JWST/NIRSpec IFU spectroscopy reveals a merging system of two galaxies (Marshall etĀ al., 2024), and multi-band JWST/NIRCam photometric observations, with the bright quasar point-spread function (PSF) subtracted, point to a combined stellar mass of the merging components less than about (Marshall etĀ al., 2024; Stone etĀ al., 2024), corresponding to even after the merger. This is further from the local relation than any other bright quasar at the end of the Epoch of Reionization yet studied (Sun etĀ al., 2025b).
If J1120+0641 is to move towards the local relation, it must accrete significant mass from surrounding galaxies, and therefore must live in a dense environment. But here, too, J1120+0641 is apparently strange: Simpson etĀ al. (2014) targeted J1120+0641 with deep HST ACS and WFC3 observations and, using a color-color selection technique to search for Lyman break galaxies, identified two galaxy candidates in their 13 arcmin2 field. This is insufficient evidence for an overdensity. In fact, the presence of only two galaxies near the quasar redshift, in a 13 arcmin2 field where galaxies are predicted from the field luminosity function, suggested that J1120+0641 may live in a significantly underdense environment. With the HST data alone, there appears to be no way for the host galaxy of J1120+0641 to gain enough mass to bring its host galaxy in line with the mass of its black hole.
However, Simpson etĀ al. (2014) had access to only three HST photometric bands, with their longest (F125W) corresponding to rest-frame nm at . Is the nondetection of a galaxy overdensity around J1120+0641 simply a result of the limited wavelength coverage of the existing HST data? If re-observed with JWST, would a galaxy overdensity reveal itself and, furthermore, reveal enough mass to grow J1120+0641ās host galaxy by the nearly two orders of magnitude necessary to bring it to the local local relation?
We designed a JWST GTO program (PID 2774) to re-observe the environment of the quasar J1120+0641. The quasar redshift places the [OĀ III] 5007 emission line within the NIRCam F405N filter. At , most galaxies should be vigorously forming stars and be bright in [OĀ III]: this narrow-band selection therefore provides a nearly complete sample of galaxies at or near the redshift of the quasar. By combining HST archival photometry and additional JWST broadband photometry with our narrow-band data, we probe the environment of J1120+0641 and examine future possibilities for this unique system.
Our HST and JWST data and their reduction are described in Section 2. We outline the candidate selection and spectral energy distribution (SED) fitting process in Section 3. Our determination of the environment of J1120+0641 is discussed in SectionĀ 4, and the implications of our results are explained in SectionĀ 5. We conclude in Section 6. Throughout this work, we use a flat cosmology with H0 = 69.6 km s-1 Mpc-1, = 0.286, and (Bennett etĀ al., 2014). All magnitudes reported are in the AB system.
2 Observations and Data Reduction
2.1 JWST NIRCam
At , the redshift of J1120+0641, the [OĀ III] 5007Ć Ā emission line is shifted to 4.05 m, into the NIRCam F405N narrow-band filter. We obtained deep observations in this filter ( = 10522.0 s) and in the wider, overlapping F410M filter ( = 5239.5 s) to search for [OĀ III]-emitting galaxies at the quasar redshift. The F360M filter ( = 5239.5 s) provides a measurement of the adjacent continuum.
We pair each of these NIRCam LW filters with a SW filter. F090W ( = 10522.0 s) straddles the Lyman break at the quasar redshift, and together with the deep archival HST F814W observations (see Section 2.2) will add valuable Lyman-break redshift constraints to exclude lower- and higher-redshift interlopers. Of particular concern is H, which enters the F405N bandpass at : these galaxies may be easily confused for [OĀ III] emitters. We therefore also observe in F150W ( = 5239.5 s) and F200W ( = 5239.5 s), which fall on either side of the Balmer break at and allow us to more easily identify and remove H contaminants.
provide additional photometric redshift constraints to exclude lower- and higher-redshift interlopers from our sample. In particular, F090W straddles the Lyman break and F814W lies blueward of it for galaxies near .
We observed the field using both NIRCam modules, with the quasar in Module B. To mitigate cosmic rays and detector artifacts, we adopted a dithering pattern, using primary dither type INTRAMODULEBOX and subpixel dither type STANDARD and the DEEP2 readout pattern.
We processed our NIRCam data using the JWST pipeline version 1.14.1 (Bushouse etĀ al., 2023), roughly following the procedures recommended in the STScI JWebbinars.111https://www.stsci.edu/jwst/science-execution/jwebbinars The pipeline parameter reference file is registered in the JWST Calibration Reference Data System (CRDS) as jwst_1235.pmap. We added a custom step to Stage 2 of the pipeline to characterize and subtract 1/f noise, a striping pattern in the images caused by the detector readout, from each frame. We produced final mosaic images in each band in Stage 3 of the pipeline by aligning and stacking all frames obtained in Stage 2, and resampled the images to a smaller pixel scale (0.0300ā³/pixel) using the drizzling algorithm in the Resample step of the pipeline. Finally, we used the Photutils Background2D function on the mosaiced and resampled images to perform a global background subtraction. We reach aperture-corrected 5 point-source depths of 0.16, 0.15, 0.16, 0.29, 1.72, and 0.42 Jy in F090W, F150W, F200W, F360M, F405N, and F410M respectively, measured in apertures containing 80% of the encircled energy for each PSF.
2.2 HST
HST images in the F814W filter were taken with the ACS wide field camera between March 27 and March 29, 2013 as part of GO-13039 (PI: Simpson). Observations were spread over 9 visits and 25 individual exposures, totaling 28,448 seconds. We re-reduced this data as follows.
Relative astrometry within a single visit was good, but small astrometric offsets existed between visits, which, if not corrected, produced artifacts in the final combined image. These were corrected by first combining the exposures taken in each visit, using astrodrizzle, then finding relative translational and rotational offsets between the combined, drizzled images using tweakreg. Alignment was done using these drizzled images to preserve the relative astrometry between exposures in a given visit, and because combining multiple images increases effective depth and reduces the impact of cosmic rays, thus improving source detection and centroiding, ultimately improving the astrometric solutions. Those solutions were propagated back to individual exposures using tweakback. Finally, the 25 astrometrically-corrected frames were combined using astrodrizzle onto north-up grids with 0.0300ā³/pixel pixels (matching the JWST data), using a Gaussian kernel and a drop size (final_pixfrac) of 0.8.
3 Analysis and Results
To generate source positions and apertures to extract photometry, we use the MAST-generated F405N segmentation image, as we are only interested in sources detected in this band. 1054 sources are identified in the F405N image.
We derive Kron radii, semi-major and semi-minor axes, and orientations for each source in the F405N segmentation image using photutils, and use these parameters to extract aperture photometry in the remaining bands. Because the F405N photometry has a lower SNR than the other bands, we explored using the F410M segmentation image to derive aperture parameters: the resulting apertures were not systematically different than those determined from the F405N image.
We use the random apertures method to calculate errors on the measured fluxes. After masking out all sources in the image, we generate 2000 random positions within the image and place the aperture and measure photometry at each. Some fraction of these generated positions invariably fall on the gap between the NIRCam modules, so we set the value of these pixels to NaN so they will not affect the noise calculated.
3.1 Candidate Identification
We require that each galaxy be present (not necessarily significantly detected) in all bands from F150W to F410M (measured flux in aperture greater than zero, within the error bars222We ensured this did not artificially limit our completeness: relaxing this requirement adds 17 additional sources, almost all of which are spurious (see next paragraph).). We expect the Lyman break to fall in the F090W filter, so we do not require that the source be present in this band (or F814W, when available). We are primarily interested in galaxies with a flux excess in the narrow band, F405N, compared to the adjacent ācontinuumā band F360M, so we make the initial color cut . 454 of the 1054 sources fit these criteria.
However, members of a larger overdensity around J1120+0641 may have a larger velocity distribution, and their [OĀ III] emission may fall outside the very narrow F405N bandpass. We also therefore include galaxies that are not elevated in F405N relative to F360M, but are elevated in the overlapping medium band F410M (). This added 69 sources to our sample, for a total of 523 candidates.
We visually inspected these 523 candidates to narrow the sample before performing SED fitting. 69 of the 523 sources are spurious: these are a mix of detector artifacts, shredded diffraction spikes of bright point sources in the image, and real galaxies only partially overlapping with the image in one or more bands (see the top panels of Figure 1 for an example). We also remove 112 galaxies obviously at very low redshift based on clearly resolved spiral structure/dust lanes (88 galaxies). Galaxies at are very unlikely to have diameters (defined as twice their effective radii) (Raikov etĀ al., 2025), so we impose a conservative cut at spatial extent greater than (103 galaxies). Finally, we perform a flux cut to remove any with fluxes inconsistent with even the most extreme galaxy (1 Jy in the UV continuum bands F090W, F150W, and/or F200W) not discarded in the visual inspection. This removes another 48 galaxies. For examples of discarded galaxies, see the lower panels of Figure 1. These low-redshift galaxies are sufficiently different from possible contaminants in our study (in SED as well as morphology) that they do not need further study.
Our visually inspected sample consists of 294 candidates elevated in F405N and/or F410M relative to F360M, which we pass to SED fitting.
3.2 Spectral Energy Distribution Fitting
For two independent determinations of the exact redshifts of our 294 candidates, we pass them through two SED fitting codes: Prospector and Bagpipes.
We adopted a modified version of the Prospector code (Johnson etĀ al., 2021), which is described in full by Lyu etĀ al. (2024). Briefly, we incorporate FSPS to model stellar emission, and model the AGN component and galaxy dust emission via semi-empirical templates. For the stellar component, we assumed a delayed-tau star formation history with a Kroupa (2001) initial mass function, the Kriek & Conroy (2013) extinction law, and the associated nebular emission components, following the default parameter priors. Given the low-mass properties of these galaxies and the dominant obscured AGN population at high- (e.g., Lyu etĀ al., 2024), we expect no significant AGN contribution to the rest-frame UV-optical SEDs of these galaxies, so the AGN component is turned off. For the galaxy dust component, we adopted the Haro 11 template without the requirement of energy balance (as we lack constraining rest-IR data).
We ran Prospector under these conditions three times for the 294 candidates identified in Section 3.1, with different tophat redshift priors: (v1), (v2), and finally (v3; the [OĀ III] line shifts into the F405N bandpass for this redshift range). Galaxies far from the quasar redshift will be worse-fit under the stricter redshift priors, yielding a larger . On the other hand, galaxies near the quasar redshift are likely to have a smaller value in v3 than in v2 or v1, and/or be fit to in the less restrictive v1 or v2 runs.
We identify 32 sources with a smaller with the most restrictive redshift prior (v3) than in v2 or v1. We also find 5 sources that have higher values in v3, but are also fit to in the lower v1 or v2 runs. Our Prospector-curated sample therefore consists of 37 galaxies total.
We also fit the 294 candidates with Bagpipes. We ran Bagpipes with a delayed-tau star formation history and Kroupa (2001) initial mass function, as with Prospector. The Kriek & Conroy (2013) dust law is not supported in Bagpipes, so we use the Salim etĀ al. (2018) law, which is parameterized similarly. We allow the maximum age of the stellar population to vary from zero to the age of the Universe at the fit redshift. The total mass of stars formed was allowed to vary from 106 to 1012 Mā, and the metallicity from 0.1% to 100% solar. Modifying the functional form of the star formation history did not significantly affect the resulting stellar mass (5-10% difference) or quality of the fit. For most sources, modifying the dust extinction law to a CF00 or Calzetti law changed the resulting stellar mass by , though in a few cases the mass changed by up to . As with Prospector, we ran Bagpipes three times with three different redshift ranges: (v1), (v2), and finally (v3). The v3 range is larger for Bagpipes than for Prospector, because Bagpipes imposes a hard limit at the edges of the range, while Prospector can fit galaxies to a redshift outside the prior range.
Bagpipes fits 24 galaxies to a v3 lower than or comparable to v2 and v1; 22 of these candidates overlap with the Prospector sample. Of the 15 galaxies fit to in Prospector but not Bagpipes, nine are only mildly elevated in F405N relative to F360M and F410M, and Bagpipes fits them to based on the position of their Lyman break. Three more are fit to , with F405N and/or F410M on H, and the remaining three are fit to . These galaxies also tend to have larger values in the Prospector fits, and tend to lie at the bright end of the Prospector sample.
We therefore proceed with the sample of 22 galaxies which both Prospector and Bagpipes robustly fit to . Their properties are summarized in Table 1. We note that the redshift range probed by the F405N filter, , is not perfectly centered at the quasar redshift (), which may introduce a sampling effect. We do not believe this makes a significant difference in our final sample, as the vast majority of our final sample have best-fit redshifts near the center of this range.
Of the 272 candidates not robustly localized to , the majority () fall into one of two categories. 118 candidates are fit by both Prospector and Bagpipes as Lyman-break galaxies at high redshift. These galaxies generally display a flat spectrum longward of F090W, and decline sharply in F090W before disappearing altogether in F814W. Another 50 galaxies are fit to by both codes. These galaxies tend to be very red across our observed bands, and are fit with older stellar populations (stellar age generally Gyr) and . These galaxies are bright in F360M, F405N, and F410M, which sit on the peak of the old stellar emission at 1.5-2 m, and the flux declines at shorter wavelengths. This decline in the bluer bands can mimic a Lyman break: a further 18 galaxies are fit to by Bagpipes but with a Lyman break () by Prospector.
3.3 Final Sample
| ID | RA (ā) | Dec (ā) | Field | aaThe uncertainties on the redshifts are small due to the narrow-band selection method; the presence of the bright F405N and the Lyman break in F090W and F814W robustly places the high-confidence candidates between . The position of the Lyman break is not as clear for the three medium-confidence candidates (see Figure 11), and they are generally not as bright in F405N relative to F360M. Their redshift errors are slightly larger, closer to . | aaThe uncertainties on the redshifts are small due to the narrow-band selection method; the presence of the bright F405N and the Lyman break in F090W and F814W robustly places the high-confidence candidates between . The position of the Lyman break is not as clear for the three medium-confidence candidates (see Figure 11), and they are generally not as bright in F405N relative to F360M. Their redshift errors are slightly larger, closer to . | /Mā) | /Mā) | |
| 300 | 169.99902 | 6.70370 | Q | 7.11 | 7.10 | |||
| 335 | 169.99734 | 6.69700 | Q | 7.10 | 7.11 | |||
| 337 | 169.99750 | 6.69704 | Q | 7.11 | 7.11 | |||
| 414 | 169.99737 | 6.68958 | Q | 7.10 | 7.12 | |||
| 427 | 169.99444 | 6.68186 | Q | 7.11 | 7.11 | |||
| 534 | 170.00581 | 6.69992 | Q | 7.11 | 7.11 | |||
| 624 | 170.00959 | 6.70138 | Q | 7.11 | 7.11 | |||
| 666 | 170.00589 | 6.68920 | Q | 7.11 | 7.11 | |||
| 703 | 170.00662 | 6.68699 | Q | 7.11 | 7.11 | |||
| 709 | 170.00678 | 6.68705 | Q | 7.07 | 7.10 | |||
| 789 | 170.01207 | 6.69086 | Q | 7.14 | 7.11 | |||
| 876 | 170.01181 | 6.68086 | Q | 7.10 | 7.11 | |||
| 941 | 170.01047 | 6.67159 | Q | 7.13 | 7.11 | |||
| 181 | 169.97607 | 6.66299 | A | 7.11 | 7.08 | |||
| 358 | 169.97943 | 6.65433 | A | 7.11 | 7.08 | |||
| 366 | 169.97875 | 6.65236 | A | 7.11 | 7.08 | |||
| 456 | 169.97744 | 6.64111 | A | 7.12 | 7.08 | |||
| 634 | 169.97744 | 6.62748 | A | 7.09 | 7.08 | |||
| 693 | 169.98564 | 6.64099 | A | 7.11 | 7.08 | |||
| 160* | 169.96941 | 6.64997 | A | 7.23 | 7.31 | |||
| 856* | 169.99940 | 6.65472 | A | 7.01 | 7.41 | |||
| 911* | 169.98969 | 6.62826 | A | 7.20 | 7.17 |
The positions of the 22 candidates are shown, along with the position of the quasar, in Figure 2, and their thumbnails and SEDs are plotted in the Appendix. Thirteen candidates lie in Module B, in the same module as the quasar; we hereafter refer to Module B as the āquasar field.ā The remaining nine candidates lie in Module A, hereafter the āadjacent fieldā (though note that an overdensity around J1120+0641 is likely to span both fields). The vast majority of these candidates are fit to a redshift between 7.07 and 7.14 by both Prospector and Bagpipes, but three of the candidates in the adjacent field are fit farther from the quasar redshift by both codes. We consider these three candidates āmedium confidence,ā and they are shown in a lighter color and with square markers in Figure 2.
Approximately a quarter of this sample displays multi-component morphologies (see cutouts in Appendix). To ensure that none of these sources were chance alignments of galaxies at two different redshifts, we performed deblended photometry on this subset of the sample using smaller, hand-placed apertures. In all cases, the two components displayed nearly identical SED shapes. It is therefore unlikely that any of these āclumpyā sources are due to projection effects.
Despite the similar model parameters input to both codes, the stellar masses returned by Prospector are larger on average than those from Bagpipes, by approximately an order of magnitude in most cases. This is mainly a consequence of our lack of data longward of rest-frame 500 nm. Bagpipes prefers to fit a very young stellar population while Prospector tends to fit an older stellar population. The two fits diverge beyond rest-frame 1 m, but without longer-wavelength data to constrain the shape of the SED we cannot rule out either possibility. Bagpipes fits the Lyman-alpha break with a damping wing, while in some cases Prospector prefers Lyman-alpha emission and additional dust extinction in the UV (see figures in the Appendix). Both codes return similar relative between sources, though the Prospector values are shifted slightly higher overall (by % on average, due to fitting the rest-UV with more dust). The best-fit star formation rates (SFRs) from both codes agree roughly with the SFR estimated from the F405N - F360M flux, i.e. from the [OĀ III] luminosity. For our purposes, the most important parameter is the stellar mass: we discuss the effect of the two codesā mass discrepancy on our results in Section 5.3.
4 The Environment of J1120+0641
4.1 HST-only Candidates
Simpson etĀ al. (2014), using HST F814W, F105W, and F125W photometry as well as Subaru and band and Spitzer/IRAC images, identified three Lyman-break candidates, which they call LBG1, LBG2, and LBG3. LBG3 lies just outside the footprint of our JWST images, but LBG1 and LBG2 are included in our data. The SEDs and thumbnails of LBG1 and LBG2 are shown in Figure 3.
LBG1 is not detected in the F405N segmentation map, meaning it is not included in our narrow band-selected sample and its redshift determination is less certain. Its observed SED shape is consistent with a galaxy (see Figure 3), and Bagpipes fits it at , primarily from the position of the Lyman break; given the width of the F090W bandpass, the uncertainty on this redshift is of order . This is in contrast to the high-confidence overdensity candidates, where the F405N excess combined with the Lyman break allows them to be fit to with much smaller redshift uncertainties. LBG1 is certainly not a low- contaminant, but without an F405N excess we cannot confirm whether it is directly associated with the quasar and simply [OĀ III]-faint, or an unrelated galaxy at slightly lower or higher redshift than the quasar. Because of its uncertain redshift, we do not include LBG1 in the candidate sample going forward.
LBG2 is robustly detected by Spitzer at 3.6 and 4.5 m. Simpson etĀ al. (2014) fit the SED of LBG2, including the Spitzer data, with EAzY, which indicated LBG2 likely lies at rather than . Our JWST data shows a steeply rising SED between F814W and F410M, unlike the rest of our candidates. LBG2 is extremely bright in the redder NIRCam bands (Jy in F405N and F410M, see top right of Figure 3), consistent with its Spitzer detections (though the IRAC1 flux of LBG1 is higher than our F360M flux). Because of this very red SED, LBG2 is selected by our narrow-band excess criterion, but is removed from our final candidate sample at the visual inspection stage because of its large flux and extended shape in the redder NIRCam filters. Bagpipes places LBG2 at .
4.2 Number Counts
We derive number counts via a Monte Carlo process. This allows us to take into account the uncertainty on source magnitudes, as each galaxyās uncertainty may shift it into a different magnitude bin. For each galaxy in the sample, we select a random value from a Gaussian centered at the objectās F150W flux, with standard deviation set to the F150W uncertainty from random apertures. We convert these resampled fluxes to apparent magnitudes, then sort the galaxies into magnitude bins. By repeating this process 1000 times, we derive a distribution of the number of galaxies in each bin. We take the mean of this distribution as the final count in the bin, and the standard deviation as its uncertainty.
The final counts in the quasar field (filled navy squares) and adjacent field (filled teal diamonds) are shown in Figure 4. For the adjacent field, we also include the counts derived when we include the three medium-confidence candidates (open teal diamonds).
Comparing these counts to Simpson etĀ al. (2014) is not trivial. Plotting the raw number of candidates detected will not properly represent the difference in sensitivity between our two samples, because the redshift range probed by our narrow-band selection method is very different from that probed by their three-band color-color technique. In theory, [OĀ III] selection with the F405N filter should probe galaxies between and , which agrees with our best-fit redshifts from SED fitting (see Table 1). A color-color selection probes a much larger redshift range, i.e. if all else is equal (depth, etc.), their technique will select more galaxies per unit area.
Therefore, to robustly compare our counts with those from Simpson etĀ al. (2014), we need to estimate the redshift range probed by their LBG selection technique, and calculate the fraction of that range that will lie within . Simpson etĀ al. (2014) verified their selection method by degrading images of the Hubble Ultra Deep Field (HUDF) to the depth of their quasar observations. They used their color-color technique to select a sample of galaxies in the HUDF, and compared it to canonical HUDF samples derived with the full filter set and depth of the HUDF. Their selection picks up five HUDF galaxies, which have robust photometric redshifts from McLure etĀ al. (2013) between and . We therefore assume that the Simpson etĀ al. (2014) color-color LBG selection technique uniformly selects galaxies between and , meaning that 1/12th of galaxies selected with their method will lie between and .
For this reason, we scale down the counts from the two Simpson etĀ al. (2014) sources by a factor of 12 and plot them in Figure 4 as teal circles; we do not perform the Monte Carlo procedure, and so do not include uncertainties on these counts.
We also plot HST counts in multiple fields from Bouwens etĀ al. (2015), the deepest of which reaches a 5 depth of mag. Like the Simpson etĀ al. (2014) data, these counts probe a much larger redshift range than our narrow-band selection technique. However, Bouwens etĀ al. (2015) provide the photometric redshift distribution of their sample (their Figure 1), which allows us to determine that approximately of their sample should lie within . We scale their data accordingly in Figure 4.
With the Simpson etĀ al. (2014) data alone, the quasar field appears slightly underdense compared to the field. However, with the addition of the JWST narrow-band selected galaxies, it is clear in Figure 4 that both the quasar field and the adjacent field are significantly overdense in galaxies. We note that we are complete only to mag F150W, which is responsible for the falling counts at fainter magnitudes. The counts in the two fields are similar down to the completeness limit, but the quasar field is slightly richer than the adjacent field. Incorporating the three medium-confidence candidates in the adjacent field does not affect this conclusion significantly.
5 Interpretation
ULAS J1120+0641 is one of the most thoroughly studied very high redshift quasars. We review the properties of the quasar and its present host galaxy in this section.
5.1 Black Hole Mass
No dynamical mass measurement is available for the SMBH in J1120+0641, but the many single-epoch measurements of its mass are listed in Table 2, based on line widths and the standard single-epoch mass determination (e.g., Vestergaard & Peterson, 2006). In general, this suite of measurements shows good agreement and indicates Mā. The single-epoch technique based on line widths can be subject to systematic errors due, for example, to outflows, so we include two other measures, both of which agree with the conclusion from the conventional single-epoch determinations.
The luminosity and temperature of the accretion disk are controlled by the black hole mass; models of the disk provide a means to estimate this mass (Calderone etĀ al., 2013). Campitiello etĀ al. (2019) used this method to determine possible masses for the black hole in J1120+0641. In TableĀ 2 we show the range of values for different viewing angles, black hole spins, and using both the relativistic model KERRBB (Li etĀ al., 2005) and a slim disk model (e.g., Sadowski etĀ al., 2009).
Recently, King (2024) suggested that many high-redshift quasars might have beamed outputs. This would raise their black hole signatures well above the true, unbeamed values, and the black hole masses calculated from observations of beamed quasars will be overestimated. We can evaluate the possibility of beamed emission from J1120+0641 from the velocities provided in Venemans etĀ al. (2017); Bosman etĀ al. (2024); and Marshall etĀ al. (2024). From the latter reference we obtain a host [OĀ III] redshift of while from the first reference we find a host redshift from [CĀ II] of (average 7.08275; weighted average 7.08496). The two broad components of H have redshifts of 7.075 and 7.0936, corresponding to outflows at -315 and 360 km s-1 respectively (Bosman etĀ al., 2024). Taking the broad-line redshift to be the average and comparing with the host galaxy redshift, the galaxy velocity relative to the broad-line region is 100 km s-1. This relatively small velocity difference is inconsistent with beaming: we therefore find no evidence that the single-epoch measurements of J1120+0641ās black hole mass are overestimated.
J1120+0641 has been detected in X-rays by Chandra and XMM-Newton (Page etĀ al., 2014; Moretti etĀ al., 2014). These two papers agree that the slope is consistent with that of typical lower-redshift X-ray quasars, i.e. that the X-ray luminosity is ānormal.ā The slope of the X-ray spectrum depends on the Eddington ratio (Moretti etĀ al., 2014), and can provide another constraint on the black hole mass. We constrain the X-ray slope using the fractional 2 - 10 keV luminosity. The softening of the slope of the spectrum reduces the fractional luminosity (or, equivalently, increases the bolometric correction ). Moretti etĀ al. (2014) report for J1120+0641, taking the bolometric luminosity of erg s-1 (see Appendix B; these errors only account for luminosity uncertainties). Page etĀ al. (2014) find of either or depending on whether the Chandra or XMM-Newton data is used to determine the 2 - 10 keV luminosity.
These estimates are widely spread, but nonetheless indicate that the luminosity is not strongly super-Eddington. The theoretical models of Pacucci & Narayan (2024) show that Eddington ratios between 1 and 2.5 lead to for quasars, much higher than indicated for J1120+0641. Their models yield lower for high Eddington ratios only if the black hole is viewed pole-on and has significant spin, but these conditions are unlikely for J1120+0641 given the lack of beaming. This implies that J1120+0641 is not accreting significantly above the Eddington limit.
Similarly, the models of Inayoshi etĀ al. (2024) (see their Figure 4) place J1120+0641 below unity Eddington ratio (i.e., ). Allowing for errors, a highly super-Eddington ratio is not plausible, although a mildly super-Eddington one is possible.
Higher super-Eddington ratios are also unlikely for J1120+0641 because they imply a black hole mass well below the broad line estimates. From these considerations, we place an upper limit on the Eddington ratio, . Since the luminosity is well determined (Appendix B), this puts a lower limit on the black hole mass as indicated in TableĀ 2. Combining all the methods, we adopt a fiducial value of Mā.
| ( Mā) | Error | Reference | |
| Broad Lines (Single-Epoch) | |||
| MgII | 2.4 | De Rosa etĀ al. (2014) | |
| CIV | 1.09 | De Rosa etĀ al. (2014) | |
| MgII & CIV | 1.35 | 0.04 | Yang etĀ al. (2021) |
| CIV | 2.40 | Farina etĀ al. (2022) | |
| H | 1.55 | 0.22 | Bosman etĀ al. (2024) |
| Pa | 0.98 | Bosman etĀ al. (2024) | |
| Pa | 0.87 | Bosman etĀ al. (2024) | |
| H | 1.4 | Marshall etĀ al. (2024) | |
| Accretion Disk Modeling | |||
| 1.3 - 7.7 | ā | Campitiello etĀ al. (2019) | |
| Eddington Ratio 2 | |||
| 1 | this work | ||
5.2 Host Galaxy Mass
Our estimate of the mass of the galaxy depends both on our understanding of the relevant observational parameters, and on the stellar initial mass function assumed. To put together a complete picture of the host galaxyās characteristics, we will discuss (1) the implications from the ULIRG-level SFR on the derived mass; (2) the mass of the stellar population seen in the UV - IR; and (3) the possible effects of the initial mass function (IMF). Given that J1120+0641 is already believed to have have a relatively low stellar mass relative to its , we will also address deriving an upper limit to the stellar mass.
5.2.1 Far infrared emission and associated SFR
The J1120+0641 host has a substantial far-infrared (far-IR) luminosity, indicating a very high SFR. Venemans etĀ al. (2020) use a rest-frame 1900 GHz (158 m) continuum measurement to derive a far-IR luminosity, assuming a modified blackbody model for the far-IR SED, with a temperature of 47K and = 1.6. They multiplied this FIR estimate by 1.41 to correct it to total infrared (TIR) luminosity. However, a modified blackbody may not be the best model at such high redshifts, and the FIR SED of low-redshift, low-metallicity Lyman continuum emitter Haro 11 should provide a more accurate model for the host (see e.g. Lyu etĀ al., 2016; De Rossi etĀ al., 2018). We therefore normalize the SED of Haro 11 to the 158 m measurement and correct to TIR by integrating, finding a correction of a factor of 1.63 (rather than 1.41) and a ULIRG-level TIR luminosity of Lā, implying an obscured SFR of 280 Mā yr-1 (Kennicutt & Evans, 2012; Murphy etĀ al., 2011; Hao etĀ al., 2011).
In general, ULIRGs absorb most of the energy in emission lines due to young stars, with only a fraction that escapes, meaning SFR estimates from optical and near-IR emission lines underestimate the total star formation. We can check for this behavior by comparing J1120+0641ās SFR from the infrared luminosity with what would be deduced from standard optical emission line estimates. The much lower SFR from H (12.5 Mā yr-1, Marshall etĀ al., 2024) seems to support this point, but because of artifacts from their PSF subtraction, the region within a radius of 02 is not included in this measurement. We expect this region to contain most of the central starburst, based on the 1 - 1.5 kpc spatial extent of the far-IR emission (Venemans etĀ al., 2017). Bosman etĀ al. (2024) measured H in an aperture of radius 03 centered on the nucleus, providing a possibility to fill in the missing area. In FigureĀ 5, we show the result of refitting their line profile, including two broad components as in their fit but adding an unresolved line. This unresolved component reduces the of the fit, based on an error estimated by repeating the fit multiple times with the narrow component set to different wavelengths. This fit is offset to the red by 200 km s-1, consistent with the result in SectionĀ 5.1 given that we are probing different regions in the galaxy. The estimated flux of the narrow H line is W m-2. A similar fit to the Pa profile, keeping the line widths and relative wavelengths (in velocity space) shows a weak detection (). The fact that neither H nor Pa show a strong narrow line supports the host galaxyās ULIRG behavior, indicating strong obscuration that suppresses optical/near-IR star formation signatures.
The SFR deduced from the H is about 40 Mā yr-1. Combining it with the Marshall etĀ al. (2024) H measurement at larger radii yields a total of Mā yr-1, less than 20% of the FIR SFR. This difference must be attributed to very strong extinction in the star-forming region, supported by the compactness and strength of the far-IR emission. Such behavior is typical of high redshift ULIRGs (Swinbank etĀ al., 2004; Takata etĀ al., 2006).
The host galaxy of J1120+0641 therefore appears very similar to lower redshift ULIRGs of similar luminosity. The typical timescale for this level of star forming activity is less than 10 Myr (e.g., Thornley etĀ al., 2000; Bekki etĀ al., 2006; Kawaguchi etĀ al., 2020). Thus the current burst of star formation can be expected to add about Mā to the host galaxy mass. These newly forming stars are hidden by the dust extinction from the rest optical and UV. We can therefore use existing HST and JWST observations of the quasar to characterize the older stellar population that dominates the mass.
5.2.2 Older, optically visible host galaxy
A nominal mass for the host galaxy stellar population of Mā is reported in Sun etĀ al. (2025b). This mass was determined by comparison with field galaxies of similar mass and redshift, as originally described in Stone etĀ al. (2024), and is based on models using a Chabrier or Kroupa initial mass function. Marshall etĀ al. (2024) estimated the mass using Prospector and a delayed-tau star formation history (with an optional late starburst). They obtained a mass of Mā. These estimates are in agreement, despite the different approaches. We adopt the average of the two, Mā, as our best estimate.
Stone etĀ al. (2024) based their estimate on values derived from modeling field galaxies, rejecting those dominanted by very young stellar populations. We show in SectionĀ 5.3 that Prospector fits such as those used by Marshall etĀ al. (2024) also return masses assuming a mature stellar population. As documented in SectionĀ 5.3, if instead the host galaxy is dominated by very young stars, then its mass could be significantly lower than this estimate.
Our mass estimate is unlikely to be āmissingā mass due to a compact stellar core behind the quasar or uncorrected dust extinction. Marshall etĀ al. (2024) found the host to be well resolved, with an effective radius reff = 0361 0001. Their images of the host display virtually identical morphologies in F115W, F200W, and F356W, with half widths at half maximum of the diffraction-limited images ranging from 002 to 006. The presence of a compact central component that could add significantly to the total galaxy flux would require a major departure from a SĆ©rsic profile and seems quite unlikely.
If the flux from the older stellar population (visible to JWST) is similarly obscured to the newly forming population, the mass will be underestimated; this may have contributed to the large positive uncertainty in the Prospector models. We can show more directly that this is unlikely to be a major issue by fitting the combined galaxy flux measurements from Stone etĀ al. (2024) and Marshall etĀ al. (2024) (reported in TableĀ 3). FigureĀ 6 shows these fits: models with a broad range of star formation histories can fit the data without invoking large extinction. In addition, there is no evidence from the photometry for extinction sufficiently strong to result in a significant underestimate of the galaxy mass.
| wavelength(m) | Flux (Jy) | error |
| 1.15 | 0.15 | 0.03 |
| 2.05 | 0.19 | 0.07 |
| 3.56 | 0.19 | 0.05 |
| 4.80 | 0.31 | 0.13 |
5.2.3 Potential total host galaxy mass
Combining the Mā of stars forming in the ULIRG process (Section 5.2.1), the Mā of older stars (Section 5.2.2), and Mā in the merging companion discovered by Marshall etĀ al. (2024), the total potential stellar mass for the host galaxy becomes Mā, assuming a local, Chabrier/Kroupa IMF for the stellar population(s).
We note that there are arguments that a top heavy IMF is more appropriate (Chon etĀ al., 2022; Steinhardt etĀ al., 2023; Bate, 2025), because the warmer cosmic background radiation during the Epoch of Reionization (e.g., 30 K at z = 10) warms the ISM, and if the ISM is metal-rich it radiates this heat away inefficiently. The warmer gas may inhibit the formation of low-mass stars and brown dwarfs. For a massive galaxy at like the J1120+0641 host, the typical metallicity is 0.1 Zā (Curti etĀ al., 2024). The models of Bate (2025) then predict a reduction of the host stellar mass of % compared to a local IMF. We proceed with the mass values for a local IMF rather than a top-heavy IMF, however, as this gives the highest masses and therefore the most conservative estimate of at .
Another contribution to the host galaxy stellar mass is its atomic and molecular gas reservoir, which does not contribute to its luminosity as observed by JWST. Values for the mass of this gas are vary significantly. Neeleman etĀ al. (2021) provide an estimate of Mā by converting the far-IR flux density from ALMA to a dust mass and assuming a dust-to-gas mass ratio. They also estimate a mass of Mā directly from the [CĀ II] luminosity. Venemans etĀ al. (2017) derive upper limits for the gas mass using two mm-wave lines with ALMA: from an upper limit on the CO(2-1) line flux; and Mā from an upper limit on the CO(7-6) line flux. All of these values are consistent with Mā, which we adopt. Only a fraction of this gas will eventually be converted into stars, governed by the star formation efficiency (SFE). Polzin etĀ al. (2024) find SFEs of a few %, independent of metallicity. Kim etĀ al. (2021) find possible SFEs of a few % to about 10%. Andalman etĀ al. (2025) run simulations for where the high gas density can favor increased SFE, and typically find values between 10% and 20%. However, estimates of the SFE are highly variable in general, depending on the assumed boundary conditions and approach. To be conservative and avoid underestimating the mass, we adopt a fiducial value of 20% for J1120+0641, in which case the gas presently in the galaxy could eventually form about Mā of stars. For the effect of other choices of SFE on the final mass, see Table 4.
Assuming that the interacting galaxy is subsumed into the host, and adding the other potential stellar contributions, we project thatānot yet accounting for accretion of any galaxies in the fieldāthe existing resources could lead to a host of about Mā, about nine times the mass of the black hole and a factor of 100 smaller than the standard Magorrian ratio (, for a predicted host mass of Mā). The effects of the uncertainty on the stellar mass and any changes to the adopted gas mass or SFE are reported in Table 4: there is approximately a factor of 2 error on this value.
| Source of uncertainty | Minimum | Adopted | Maximum | Unit | Reference(s) |
| CURRENT HOST GALAXY | |||||
| Stars | |||||
| Stellar mass at | 0.57 | 0.76 | 1.05 | Mā | M24 |
| Gas/Future Stars | |||||
| Stellar mass forming in current starburst | 0.05 | 0.10 | 0.20 | Mā | this work |
| Gas mass at | 0.50 | 2.00 | 3.00 | Mā | N21, V17 |
| Star formation efficiency to | 0.10 | 0.20 | 0.40 | An25, K21, P24 | |
| = Stellar mass formed by | 0.05 | 0.40 | 1.20 | Mā | |
| Stellar Mass at | 0.67 | 1.26 | 2.45 | Mā | |
| SATELLITE GALAXIES (QUASAR FIELD ONLY) | |||||
| Stars | |||||
| Stellar mass detected | 0.20 | 0.71 | 1.21 | Mā | this work |
| + Stellar mass missed (incompleteness) | 0.18 | 0.35 | 0.53 | Mā | |
| + Stellar mass missed (no [OĀ III]) | 0.19 | 0.52 | 0.85 | Mā | |
| = Total stellar mass () | 0.57 | 1.58 | 2.59 | Mā | |
| Gas/Future Stars | |||||
| at | 5 | 10 | 20 | Al25, H22 | |
| Gas mass at | 2.85 | 15.8 | 51.8 | Mā | |
| Star formation efficiency to | 0.1 | 0.2 | 0.4 | An25, K21, P24 | |
| = Stellar mass formed by | 0.29 | 3.15 | 20.7 | Mā | |
| Stellar Mass at | 0.86 | 4.73 | 23.3 | Mā | |
| TOTAL | |||||
| Host Stellar Mass at : | 1.53 | 5.99 | 25.8 | Mā | |
5.3 The Future of ULAS J1120+0641
The host galaxy of J1120+0641 can grow via mergers with its surrounding galaxies between and . Here, we estimate the total mass available to be accreted and to form stars.
As discussed in Section 3.3, the two SED fitting codes we use return very different masses for the surrounding galaxies due to our lack of wavelength coverage in the (rest-frame) red and near-IR. Bagpipes, which prefers to fit extremely young stellar populations (mass-weighted age Myr) for almost all sources, returns masses between and Mā for the high-confidence candidates. Prospector, which fits more evolved stellar populations with smaller emission-line contributions, yields to Mā for the same objects, approximately an order of magnitude higher in most cases. The masses from Prospector at lower redshift and on relatively old galaxies have been shown to agree well on average with dynamical masses (Cappellari, 2023). Given comparable initial conditions, Prospector and Bagpipes return very similar mass estimates (Pacifici etĀ al., 2023). That is, the differences in mass estimates do not reflect uncertainties in the modeling programs. Indeed, we find that if Bagpipes is forced to fit an older stellar population, it returns very similar results to Prospector (albeit with a larger -squared for the fit). The Prospector masses are traceable to dynamical masses if the galaxy masses are dominated by relatively old stellar populations and the Bagpipes masses represent the situation if very young stars are dominant.
As such, to estimate the total mass of our candidate [OĀ III] emitters and its uncertainty, we adopt the Bagpipes value as a lower limit on the stellar mass and the Prospector value as an upper limit, and adopt the median value going forward. Table 4 reports the uncertainty that this choice, and all others described below, have on the total inferred mass.
Under this choice, the narrow-band candidates across both fields have a total stellar mass of approximately Mā ( Mā if the medium-confidence sources, two of which are fit to dubiously high masses, are included). Mā of galaxies lies in the quasar field. All galaxies in the quasar field are high-confidence sources (see Figure 2).
It is clear from Figure 4 that the adjacent field away from the quasar is not a true blank field. The similar density of galaxies observed in the quasar and adjacent fields implies that both fields trace an overdensity spanning pMpc. However, even if the overdensity extends over a very large area, not all galaxies in the overdensity are likely to merge with the quasar host in the finite time between the quasarās observed redshift and the present day. Chamberlain etĀ al. (2024) examine Illustris TNG100 galaxy pairs with similar stellar masses to our sources, and find that galaxy pairs that merge by generally have separations kpc, even at the highest redshift they consider (). The candidates in the quasar field have projected distances kpc from the quasar, while the candidates in the adjacent field have projected distances kpc from the quasar. Any galaxies off the edge of our JWST data will also have projected distances kpc. It is therefore realistic and conservative to assume that only the sources in the quasar field could merge with the quasar host to increase its mass: Mā of stars.
However, some galaxies around the quasarāand therefore some mass potentially available to accrete onto the hostāmay be missing from our sample. The effects of incompleteness are visible at , and any galaxies without bright [OĀ III] will be missed by our narrow-band selection.
To assess the effects of incompleteness, we extrapolate our counts assuming the true distribution of galaxies follows the shape of the Bouwens etĀ al. (2015) counts to the faint end (30th mag). We scale the Bouwens etĀ al. (2015) HST counts to match our JWST counts at , which yields and at and respectively. This predicts that galaxies with and galaxies with should be present in the 4.4 arcmin2 quasar field. We detect 4 and 2 galaxies at these magnitude ranges, respectively, and use their masses to estimate the mass of āmissingā galaxies, yielding Mā missed due to incompleteness: the total stellar mass in the quasar field is then Mā.
We expect the majority of sources at to be highly star-forming and therefore bright in [OĀ III], in line with the bulk of high-redshift galaxies detected with JWST. However, it is important to understand what fraction of galaxies our selection technique will miss.
The JADES survey (Eisenstein etĀ al., 2023) covers a much larger area than our observations (45 arcmin2) to great depth, detecting hundreds of galaxies and assigning robust photometric redshifts. JADES does not observe in F405N or F360M, two of the primary bands used in our sample selection, but does use F410M and F200W. Comparing the ratio of these two bands allows us to estimate the fraction of galaxies that lack a bright [OĀ III] line. In principle, any galaxy with F410M brighter than F200W will be selected by our criteria described in Section 3.1. In practice, however, galaxies with brighter [OĀ III] are more easily fit to the quasar redshift by SED fitting codes. For this reason, the minimum observed color in our sample is 0.7.
We examine the colors of all JADES galaxies which are a) detected in both F200W and F410M, b) are fit to photometric redshifts , and c) have F200W magnitudes between 26.5 and 30, to match our sample. 67.2% of these JADES galaxies have colors consistent with our sample: this indicates that approximately 32.8% of galaxies do not have bright enough [OĀ III] in the bands to be effectively fit by the SED fitting codes. We note that the very bright sources in JADES (), almost none of which have a strong F410M excess (see Figure 7, are mostly misclassified stars. The population of galaxies without bright [OĀ III] may not follow the same spatial distribution or mass distribution as the [OĀ III] emitters: Champagne etĀ al. (2025b) examined the environment of the quasar J0305-3150 and found a correlation between [OĀ III] equivalent width and distance from the central quasar. They also found that the Lyman-break galaxies near the quasar, without [OĀ III] emission, were universally older, less star-forming, and more massive. However, without more knowledge of the population of galaxies around J1120+0641 without bright [OĀ III], we cannot robustly correct for any difference in the populations. We therefore assume that galaxies without strong [OĀ III] follow the same mass distribution as the detected sources, and correct the detected stellar masses up assuming 32.8% of the mass is missed. This adds another Mā, and brings the total stellar mass in the quasar field to Mā.
With a robust estimate of the total stellar mass in the satellite galaxies, we finally need to account for the gas present in the satellite galaxies, some fraction of which will form stars. While no comprehensive surveys from e.g. ALMA exist at the high redshift and very low mass of our sample galaxies, galaxies of slightly higher mass at the same redshift tend to have gas masses approximately an order of magnitude larger than their stellar masses (Algera etĀ al., 2025; de Graaff etĀ al., 2024; Heintz etĀ al., 2022). Assuming a gas-to-stars ratio of 10 and an SFE of 20% (see references in Section 5.2.3), the total stellar mass of Mā in the surrounding galaxies implies that an additional Mā of stars can form by . For the effect of changing the adopted gas-to-stars ratio and SFE on the final mass, see 4. Between the detected galaxies, the galaxies missed due to incompleteness, and their gas, we therefore conclude that approximately Mā of stars could possibly be added to the host galaxyās mass between its observed redshift and through mergers. This value is highly uncertain, dependent on the assumed gas-to-stars ratio and star formation efficiency: this mass from āfuture starsā dominates the uncertainty on our final host galaxy mass (see Table 4). To see the effect that different assumptions have on the derived total mass, see Table 4.
Combining the mass of the surrounding galaxies with that of the host gives a total mass of Mā. Assuming that no further black hole growth occurs, then the host of J1120+0641 will grow to , or 2.5%, at : still about a factor of 25 greater than the typical value for low-redshift galaxies of similar stellar mass (, Greene etĀ al., 2020). Only as we approach the upper end of our uncertainty rangeāstill assuming no further black hole growth, and adopting the large Prospector masses for the satellites as well as large gas fractions and very efficient star formationādo we achieve a final stellar mass of Mā, for (still on the lower end of the expected stellar mass given the mass of J1120+0641ās black hole). The range in that J1120+0641 could reach is shown in Figure 8.
Galaxies with are essentially unheard of in the local Universe. Of the sources in Greene etĀ al. (2020) with similar black hole masses to J1120+0641, none have stellar masses below Mā. The few existing examples of such extreme systems are unique cases: NGC 1277 was originally thought to have a black hole mass of Mā and of 6.3% (van den Bosch etĀ al., 2012), but later modeling (Emsellem, 2013; Graham etĀ al., 2016) found that the dynamics of the bulge could also be explained by a bar viewed end-on, and revised the black hole mass down, closer to solar masses (). NGC 4342 has a black hole mass of Mā (, Cretton & van den Bosch, 1999), but there is evidence that it is being tidally stripped by its neighbor, NGC 4365 (Blom etĀ al., 2014), reducing its stellar mass. M32 is another classic example of this phenomenon, tidally stripped by interactions with M31. There appears to exist no example of a galaxy with similar to that implied for J1120+0641 by the total mass around it.
It is therefore clear that the galaxies visible in the immediate surroundings of J1120+0641 are likely not sufficient to transport it onto the local relation. Sub-mm data to constrain the gas mass of these satellites, or typical low-mass galaxies at , could reduce our uncertainty significantly. If the host is to grow further, it must either accrete significantly more gas from the surrounding intergalactic medium or undergo a merger of its dark matter halo with another halo not visible in our observed field. Predicting this is, of course, impossible. However, if there are many other examples where this process is invoked, it would imply that the Magorrian relation is achieved in those cases stochastically, and not in a repeatable fashion.
Itās important to note that the presence of even a single massive NIRCam-dark galaxy in the vicinity of J1120+0641, similar to those discovered around the quasars J0305-3150 and J1526-2050 by Sun etĀ al. (2025a) at , could significantly increase the mass available to accrete onto the quasar host. No [CĀ II] imaging of J1120+0641 currently exists over a large enough area to confirm or refute this possibility.
It may be that the fate of J1120+0641 is in fact to evolve into a SMBH with very little host galaxy. Then the absence of known similar cases in the local Universe may be due to selection effects (but note the example of such a system at intermediate redshift presented by Schramm etĀ al., 2019). At the end of the current quasar phase, the SMBH would likely not be elevated to quasar luminosity again, due to the apparent dearth of gas around the galaxy to accrete, and therefore evolve quietly to low redshift. A quiescent SMBH in a low luminosity galaxy might escape detection altogether if it were sufficiently distant.
5.4 The Past of ULAS J1120+0641
It is emerging from JWST observations that, at high redshift, there are many examples of black holes overmassive relative to their host galaxies (e.g., Harikane etĀ al., 2023; Pacucci etĀ al., 2023; Ćbler etĀ al., 2023; Kokorev etĀ al., 2023; Goulding etĀ al., 2023; Furtak etĀ al., 2024; Juodžbalis etĀ al., 2024; Natarajan etĀ al., 2024). For the most massive cases (log(), the available sample sizes are inadequate to show whether this behavior is typical or is confined to a subsample that is being found preferentially (Sun etĀ al., 2025b). However, there are individual cases that stand out; among these, J1120+0641, with , is unique in the amount and quality of data to evaluate this behavior. As a result, concerns about the validity of the mass measurements of the black hole (Lupi etĀ al., 2024b; King, 2024) (and of the host galaxy) are greatly reduced and we can confidently explore the implications for its past evolution.
The large black hole mass of J1120+0641 only 750 Gyr after the Big Bang indicates a massive seed black hole. To reach this mass via constant Eddington accretion requires a seed mass of Mā if the accretion begins at and 4400 Mā if the accretion begins at . These seed masses resemble those predicted by direct-collapse theory, where a pristine gas cloud is prevented from fragmenting by Lyman-Werner radiation from nearby stars (e.g., Agarwal etĀ al., 2013; Chiaki etĀ al., 2023), though a number of physical challenges to this formation pathway exist (Bhowmick etĀ al., 2022a, b; Dunn etĀ al., 2018; Smith etĀ al., 2017).
If this formation pathway is possible, J1120+0641 may be an example of the evolution of an Overly Massive Black hole Galaxy (OMBG, Agarwal etĀ al., 2013; Natarajan etĀ al., 2017; Scoggins etĀ al., 2023; Scoggins & Haiman, 2024), where a massive black hole forms via direct collapse in a satellite subhalo of a star-forming halo which provides the Lyman-Werner radiation flux. These systems are predicted to have very large (greater than 1) until their host halo merges with the nearby larger halo, at which point the ratio is sub-unity, but still much greater than observed in the local Universe.
Alternatively, with episodes of super-Eddington accretion a large black hole mass can be achieved with a lower-mass seed and with a different accretion history (see e.g. Hu etĀ al., 2022; Lupi etĀ al., 2024a; Trinca etĀ al., 2024). In the detailed simulation by Lupi etĀ al. (2024a), the super-Eddington phase is delayed since supernova feedback agitates the interstellar gas and prevents inflow and accretion. However, once the galaxy grows in mass to 1010 Mā, accretion can occur very rapidly and grow the black hole by orders of magnitude. A similar result is obtained in the independent simulation by HuÅ”ko etĀ al. (2025). This hypothesis would make cases like ULAS J1120 rare, particularly at higher redshifts before galaxies have grown to sufficient mass. In comparison, Trinca etĀ al. (2024) propose that super-Eddington acretion is triggered by large scale dynamical events such as major mergers that can occur over an extended period of time. In this type of model there are bursty episodes of super-Eddington accretion throughout the early life of the host galaxy.
6 Summary and Conclusions
Using new, deep JWST/NIRCam imaging, we searched for [OĀ III] emitters around the quasar ULAS J1120+0641 to constrain the density of the quasarās environment. We use a narrow-band selection technique combined with modeling of the broadband UV/optical SED to obtain robust redshifts and masses.
ā¢
We place 19 galaxies at the quasar redshift with high confidence (3 at medium confidence), 13 of which lie in the quasar field, within a projected distance of 350 kpc. The remaining 6 high-confidence candidates and the 3 medium-confidence candidates lie in the adjacent field, with projected distances from the quasar kpc.
ā¢
We determine that two of the Lyman-break galaxies detected in the quasar field by Simpson etĀ al. (2014) are not directly associated with the quasar. LBG1 lies at and LBG2 lies at .
ā¢
J1120+0641 lies in an overdensity of galaxies spanning more than 1 proper Mpc. Both the quasar field and adjacent field are significantly overdense in sources down to our completeness limit, most of which could not be identified with HST alone.
ā¢
We combine multiple observations of the quasar to better constrain its properties, arriving at a current black hole mass Mā. Masses derived from single-epoch spectroscopy, modeling of the accretion disk, and Eddington ratio arguments are in rough agreement. This high mass less than 1 Gyr after the Big Bang likely requires a massive black hole seed and/or significant episodes of super-Eddington accretion
ā¢
We constrain the properties of the host galaxy by combining data from Marshall etĀ al. (2024), Yue etĀ al. (2024), Stone etĀ al. (2024), and Bosman etĀ al. (2024). The host resembles a ULIRG, with high levels of dust-obscured star formation in the galaxyās center, likely ignited by an ongoing merger. The two merging components of the quasar host have a total stellar mass of Mā, with an ongoing starburst, and gas supply with the potential to increase this mass.
ā¢
After accounting for incompleteness and selection effects, the stellar mass in the galaxies within kpc (projected) of the quasar is Mā. If all the galaxies in the overdensity within kpc of the quasar merge with the host galaxy by and form additional stars, the host of J1120+0641 will reach a final stellar mass of Mā (with a large uncertainty dominated by the error in the gas masses). If the black hole grows no further, this corresponds to a black hole massāstellar mass ratio of %: significantly larger than any known galaxy in the local Universe that has not lost significant stellar mass to tidal stripping. Any further growth of the host galaxy will be stochastic, through e.g. unexpectedly high accretion of gas from the intergalactic medium and/or a merger with another halo.
[…]
Appendix A Galaxy images
Appendix B Quasar luminosity
Our determination of the luminosity of J1120+0641 is centered on FigureĀ 12. The presentation shows relative luminosity for different wavelengths directly. The luminosity largely emerges between 0.05 and 30 m. The spectra and photometry cover a significant fraction of this wavelength range, except for m. The extinction to the BLR is minimal (Bosman etĀ al., 2024), so the SED short of 0.1 m should be reasonably accurate. We therefore estimate the luminosity for the normal and WDD cases Lyu etĀ al. (2017) by integrating the output for the appropriate SED, finding Lā and Lā, respectively for ānormalā and WDD. Extrapolating from Figure 12 in Lyu etĀ al. (2017), it appears that there is about a 30% probability that the WDD template is appropriate. That is, the most likely luminosity is Lā, or ergs s-1. .